# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 21.1.2 ESRGAN (Enhanced Super-Resolution GAN)
Enhanced Super-Resolution GAN (ESRGAN) is an improved version of SRGAN that addresses some limitations and introduces more sophisticated techniques to generate even more realistic and detailed images [206]. ESRGAN improves on SRGAN in several ways:
@@
1class RRDBBlock(nn.Module):
@@
4 self.conv1 = nn.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
5 self.conv2 = nn.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
6 self.conv3 = nn.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
7 self.relu = nn.PReLU()
@@ 21.1.3 VDSR (Very Deep Super-Resolution Network)
VDSR (Very Deep Super-Resolution Network) is another deep learning model designed for image super-resolution. Unlike SRGAN, which uses adversarial loss, VDSR is a fully convolutional network with many layers that learns residual mappings. By training on residuals (i.e., the difference between low-resolution and high-resolution images), VDSR is able to upscale images with high precision [85].
@@
1class VDSR(nn.Module):
@@
4 self.conv1 = nn.Conv2d(1, 64, kernel_size=3, padding=1)
5 self.relu = nn.ReLU(inplace=True)
6 self.convs = nn.Sequential(*[
7 nn.Conv2d(64, 64, kernel_size=3, padding=1),
8 nn.ReLU(inplace=True)
@@
10 self.conv_last = nn.Conv2d(64, 1, kernel_size=3, padding=1)
@@ 21.1.4 EDSR (Enhanced Deep Super-Resolution Network)
Enhanced Deep Super-Resolution Network (EDSR) is an optimized version of VDSR designed to improve both accuracy and efficiency. EDSR removes batch normalization layers, which can introduce artifacts in the image, and focuses on a simpler architecture while maintaining deep layers to handle more complex super-resolution tasks [109].
@@
1class EDSR_ResidualBlock(nn.Module):
@@
4 self.conv1 = nn.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
5 self.relu = nn.ReLU(inplace=True)
6 self.conv2 = nn.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
@@ 21.3.1 DeBlurGAN
6 class DeBlurGAN_Generator(nn.Module):
@@
9 self.encoder = nn.Sequential(
10 nn.Conv2d(3, 64, kernel_size=7, stride=1, padding=3),
11 nn.ReLU(inplace=True),
12 nn.Conv2d(64, 128, kernel_size=5, stride=2, padding=2),
13 nn.ReLU(inplace=True)
@@
15 self.decoder = nn.Sequential(
16 nn.ConvTranspose2d(128, 64, kernel_size=4, stride=2, padding=1),
17 nn.ReLU(inplace=True),
18 nn.Conv2d(64, 3, kernel_size=7, stride=1, padding=3),
19 nn.Tanh() # Output between -1 and 1
@@
28 class DeBlurGAN_Discriminator(nn.Module):
@@
31 self.network = nn.Sequential(
32 nn.Conv2d(3, 64, kernel_size=4, stride=2, padding=1),
33 nn.LeakyReLU(0.2, inplace=True),
34 nn.Conv2d(64, 128, kernel_size=4, stride=2, padding=1),
35 nn.LeakyReLU(0.2, inplace=True),
36 nn.Conv2d(128, 1, kernel_size=4, stride=1, padding=1),
37 nn.Sigmoid()
@@
48 adversarial_loss = nn.BCELoss()
49 reconstruction_loss = nn.L1Loss() # For image reconstruction
@@ 21.3.2 CNN-based Deblurring
1 class CNN_Deblurring(nn.Module):
@@
4 self.network = nn.Sequential(
5 nn.Conv2d(3, 64, kernel_size=3, stride=1, padding=1),
6 nn.ReLU(inplace=True),
7 nn.Conv2d(64, 64, kernel_size=3, stride=1, padding=1),
8 nn.ReLU(inplace=True),
9 nn.Conv2d(64, 3, kernel_size=3, stride=1, padding=1) # Output layer
@@ 21.3.3 Dynamic Deblurring
1 class DynamicDeblurring(nn.Module):
@@
5 self.motion_estimator = nn.Sequential(
6 nn.Conv2d(3, 64, kernel_size=3, stride=1, padding=1),
7 nn.ReLU(inplace=True),
8 nn.Conv2d(64, 128, kernel_size=3, stride=1, padding=1),
9 nn.ReLU(inplace=True)
@@
12 self.deblurring = nn.Sequential(
13 nn.Conv2d(128, 64, kernel_size=3, stride=1, padding=1),
14 nn.ReLU(inplace=True),
15 nn.Conv2d(64, 3, kernel_size=3, stride=1, padding=1)
@@
Diffusion models represent a powerful new paradigm for image generation, using a probabilistic framework to iteratively add and remove noise from images. They have gained significant attention in the field of generative modeling due to their ability to produce high-quality images with fine details [40]. This chapter introduces the fundamental concepts of diffusion models, explains how they operate through forward and reverse diffusion processes, and explores their applications in tasks such as image generation and denoising.
@@ 22.1 Fundamentals of Diffusion Models
Diffusion models are generative models that transform an image by progressively adding noise to it in a forward process and then reconstruct the image by reversing the process through denoising. Unlike adversarial models like GANs, diffusion models generate images by explicitly modeling the noise and using a probabilistic framework to remove it step by step [73].
@@ 22.1.1 Forward and Reverse Diffusion Process
The forward and reverse diffusion processes are at the core of how diffusion models generate images [215].
@@
16class SimpleDiffusionModel(nn.Module):
@@
19 self.denoise_net = nn.Sequential(
20 nn.Conv2d(3, 64, kernel_size=3, padding=1),
21 nn.ReLU(),
22 nn.Conv2d(64, 64, kernel_size=3, padding=1),
23 nn.ReLU(),
24 nn.Conv2d(64, 3, kernel_size=3, padding=1) # Predict denoised output
@@ 22.1.2 Probabilistic Diffusion Models
Probabilistic diffusion models are built on a well-defined probabilistic framework. They explicitly model the forward process of adding noise to an image and the reverse process of denoising the image. The model learns to approximate the reverse distribution $p(x_{t-1}|x_{t})$ using neural networks, which allows it to generate high-quality images [73].
@@ 22.2 Applications of Diffusion Models in Image Generation
Diffusion models have been successfully applied in various image generation tasks, particularly for generating high-resolution and detailed images. Two of the most popular types of diffusion models are Denoising Diffusion Probabilistic Models (DDPM) and Latent Diffusion Models (LDM), each offering unique strengths in image generation [215].
@@ 22.2.1 Denoising Diffusion Probabilistic Models (DDPM)
Denoising Diffusion Probabilistic Models (DDPM) are a class of generative models that achieve state-of-the-art results in image generation by iteratively removing noise from noisy inputs. The process is highly structured, with each time step focusing on reducing a small amount of noise, making DDPMs capable of generating high-quality images [73].
@@
1class DDPM(nn.Module):
@@
5 self.denoise_net = nn.Sequential(
6 nn.Conv2d(3, 64, kernel_size=3, padding=1),
7 nn.ReLU(),
8 nn.Conv2d(64, 64, kernel_size=3, padding=1),
9 nn.ReLU(),
10 nn.Conv2d(64, 3, kernel_size=3, padding=1) # Predict denoised output
@@ 22.2.2 Latent Diffusion Models (LDM)
  • High-quality generation: Despite operating in a lower-dimensional space, LDMs can still generate high-quality images by learning a mapping between the latent space and the pixel space [169].
@@
1class LatentDiffusionModel(nn.Module):
@@
4 self.encoder = nn.Conv2d(3, latent_dim, kernel_size=3, padding=1)
5 self.decoder = nn.ConvTranspose2d(latent_dim, 3, kernel_size=3, padding=1)
@@ 22.4.1 Image Denoising
6 class DiffusionModel(nn.Module):
@@
9 self.network = nn.Sequential(
10 nn.Conv2d(3, 64, kernel_size=3, padding=1),
11 nn.ReLU(inplace=True),
12 nn.Conv2d(64, 128, kernel_size=3, padding=1),
13 nn.ReLU(inplace=True),
14 nn.Conv2d(128, 3, kernel_size=3, padding=1)
@@ 22.4.2 Image Inpainting
1 class InpaintingDiffusionModel(nn.Module):
@@
4 self.network = nn.Sequential(
5 nn.Conv2d(3, 64, kernel_size=3, padding=1),
6 nn.ReLU(inplace=True),
7 nn.Conv2d(64, 128, kernel_size=3, padding=1),
8 nn.ReLU(inplace=True),
9 nn.Conv2d(128, 3, kernel_size=3, padding=1)
@@ 22.4.3 Image Super-Resolution
1 class SuperResolutionDiffusionModel(nn.Module):
@@
4 self.network = nn.Sequential(
5 nn.Conv2d(3, 64, kernel_size=3, padding=1),
6 nn.ReLU(inplace=True),
7 nn.Conv2d(64, 128, kernel_size=3, padding=1),
8 nn.ReLU(inplace=True),
9 nn.ConvTranspose2d(128, 3, kernel_size=3, stride=2, padding=1, output_padding=1) # Upsample
@@
20 return nn.functional.interpolate(image, scale_factor=scale_factor, mode=’bilinear’)
@@
Deep learning has revolutionized the field of image processing, enabling powerful methods for generating, enhancing, and understanding images. From creative image synthesis to enhancing resolution and clarity, deep learning models such as GANs, diffusion models, and others have opened new avenues for solving complex image-related tasks. This chapter explores advanced applications of deep learning in image generation, enhancement, and interpretation, providing a detailed explanation of the underlying techniques and their practical implications [82, 28, 161].
@@ 23.1.1 Deep Learning-based Image Generation
Deep learning-based image generation techniques involve training models to learn the statistical patterns in a large set of images. Once trained, these models can produce novel images that resemble the patterns seen during training. Two of the most prominent approaches for image generation are Generative Adversarial Networks (GANs) and Diffusion Models. Both methods use neural networks to generate images, but their underlying mechanisms differ significantly [152].
@@
: GANs are a class of deep learning models that consist of two neural networks: a generator and a discriminator, which compete in a game-like scenario. The generator tries to create realistic images, while the discriminator evaluates the images and distinguishes between real and generated ones. Over time, the generator learns to produce highly realistic images, while the discriminator becomes better at identifying fake images. This adversarial training process leads to a generator capable of creating visually appealing and detailed images [157].
@@
5class GANGenerator(nn.Module):
@@
8 self.fc = nn.Sequential(
9 nn.Linear(noise_dim, 256),
10 nn.ReLU(),
11 nn.Linear(256, 512),
12 nn.ReLU(),
13 nn.Linear(512, img_dim),
14 nn.Tanh() # Output in range [-1, 1]
@@
21class GANDiscriminator(nn.Module):
@@
24 self.fc = nn.Sequential(
25 nn.Linear(img_dim, 512),
26 nn.LeakyReLU(0.2),
27 nn.Linear(512, 256),
28 nn.LeakyReLU(0.2),
29 nn.Linear(256, 1),
30 nn.Sigmoid() # Output a probability
@@
  • The generator takes a random noise vector $z$ as input and transforms it into a high-dimensional image. The output is a flattened image of size 28x28 pixels, and the Tanh activation function ensures the pixel values are in the range $[-1,1]$.
@@
  • Image-to-Image Translation: GANs are used to convert images from one domain to another, such as turning sketches into fully colored images [218].
  • Style Transfer: GANs can blend the artistic style of one image with the content of another to create new, stylized images [54].
  • Text-to-Image Generation: GANs can generate images from text descriptions, producing realistic visuals based on the input text [107].
@@
: Diffusion models represent a different approach to image generation. Unlike GANs, which use a generator-discriminator framework, diffusion models operate by progressively transforming a simple, noisy image into a high-quality image through multiple denoising steps. These models add noise to an image in a forward process and then use a learned reverse process to remove the noise step by step, eventually recovering a clear image [82].
@@
9class ReverseDiffusionModel(nn.Module):
@@
12 self.denoise_net = nn.Sequential(
13 nn.Conv2d(3, 64, kernel_size=3, padding=1),
14 nn.ReLU(),
15 nn.Conv2d(64, 64, kernel_size=3, padding=1),
16 nn.ReLU(),
17 nn.Conv2d(64, 3, kernel_size=3, padding=1) # Output denoised image
@@
Diffusion models have become particularly popular due to their stability during training and ability to generate high-resolution images [215]. They are well-suited for applications that require fine-grained control over image generation, such as:

### v2

@@ 21.1.2 ESRGAN (Enhanced Super-Resolution GAN)
Enhanced Super-Resolution GAN (ESRGAN) is an improved version of SRGAN that addresses some limitations and introduces more sophisticated techniques to generate even more realistic and detailed images [wang2018esrgan]. ESRGAN improves on SRGAN in several ways:
@@
1class RRDBBlock($n$.Module):
@@
4 self.conv1 = $n$.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
5 self.conv2 = $n$.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
6 self.conv3 = $n$.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
7 self.relu = $n$.PReLU()
@@ 21.1.3 VDSR (Very Deep Super-Resolution Network)
VDSR (Very Deep Super-Resolution Network) is another deep learning model designed for image super-resolution. Unlike SRGAN, which uses adversarial loss, VDSR is a fully convolutional network with many layers that learns residual mappings. By training on residuals (i.e., the difference between low-resolution and high-resolution images), VDSR is able to upscale images with high precision [kim2016accurate].
@@
1class VDSR($n$.Module):
@@
4 self.conv1 = $n$.Conv2d(1, 64, kernel_size=3, padding=1)
5 self.relu = $n$.ReLU(inplace=True)
6 self.convs = $n$.Sequential(*[
7 $n$.Conv2d(64, 64, kernel_size=3, padding=1),
8 $n$.ReLU(inplace=True)
@@
10 self.conv_last = $n$.Conv2d(64, 1, kernel_size=3, padding=1)
@@ 21.1.4 EDSR (Enhanced Deep Super-Resolution Network)
Enhanced Deep Super-Resolution Network (EDSR) is an optimized version of VDSR designed to improve both accuracy and efficiency. EDSR removes batch normalization layers, which can introduce artifacts in the image, and focuses on a simpler architecture while maintaining deep layers to handle more complex super-resolution tasks [lim2017enhanced].
@@
1class EDSR_ResidualBlock($n$.Module):
@@
4 self.conv1 = $n$.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
5 self.relu = $n$.ReLU(inplace=True)
6 self.conv2 = $n$.Conv2d(num_channels, num_channels, kernel_size=3, padding=1)
@@ 21.3.1 DeBlurGAN
6 class DeBlurGAN_Generator($n$.Module):
@@
9 self.encoder = $n$.Sequential(
10 $n$.Conv2d(3, 64, kernel_size=7, stride=1, padding=3),
11 $n$.ReLU(inplace=True),
12 $n$.Conv2d(64, 128, kernel_size=5, stride=2, padding=2),
13 $n$.ReLU(inplace=True)
@@
15 self.decoder = $n$.Sequential(
16 $n$.ConvTranspose2d(128, 64, kernel_size=4, stride=2, padding=1),
17 $n$.ReLU(inplace=True),
18 $n$.Conv2d(64, 3, kernel_size=7, stride=1, padding=3),
19 $n$.Tanh() # Output between -1 and 1
@@
28 class DeBlurGAN_Discriminator($n$.Module):
@@
31 self.network = $n$.Sequential(
32 $n$.Conv2d(3, 64, kernel_size=4, stride=2, padding=1),
33 $n$.LeakyReLU(0.2, inplace=True),
34 $n$.Conv2d(64, 128, kernel_size=4, stride=2, padding=1),
35 $n$.LeakyReLU(0.2, inplace=True),
36 $n$.Conv2d(128, 1, kernel_size=4, stride=1, padding=1),
37 $n$.Sigmoid()
@@
48 adversarial_loss = $n$.BCELoss()
49 reconstruction_loss = $n$.L1Loss() # For image reconstruction
@@ 21.3.2 CNN-based Deblurring
1 class CNN_Deblurring($n$.Module):
@@
4 self.network = $n$.Sequential(
5 $n$.Conv2d(3, 64, kernel_size=3, stride=1, padding=1),
6 $n$.ReLU(inplace=True),
7 $n$.Conv2d(64, 64, kernel_size=3, stride=1, padding=1),
8 $n$.ReLU(inplace=True),
9 $n$.Conv2d(64, 3, kernel_size=3, stride=1, padding=1) # Output layer
@@ 21.3.3 Dynamic Deblurring
1 class DynamicDeblurring($n$.Module):
@@
5 self.motion_estimator = $n$.Sequential(
6 $n$.Conv2d(3, 64, kernel_size=3, stride=1, padding=1),
7 $n$.ReLU(inplace=True),
8 $n$.Conv2d(64, 128, kernel_size=3, stride=1, padding=1),
9 $n$.ReLU(inplace=True)
@@
12 self.deblurring = $n$.Sequential(
13 $n$.Conv2d(128, 64, kernel_size=3, stride=1, padding=1),
14 $n$.ReLU(inplace=True),
15 $n$.Conv2d(64, 3, kernel_size=3, stride=1, padding=1)
@@
Diffusion models represent a powerful new paradigm for image generation, using a probabilistic framework to iteratively add and remove noise from images. They have gained significant attention in the field of generative modeling due to their ability to produce high-quality images with fine details [croitoru2023diffusion]. This chapter introduces the fundamental concepts of diffusion models, explains how they operate through forward and reverse diffusion processes, and explores their applications in tasks such as image generation and denoising.
@@ 22.1 Fundamentals of Diffusion Models
Diffusion models are generative models that transform an image by progressively adding noise to it in a forward process and then reconstruct the image by reversing the process through denoising. Unlike adversarial models like GANs, diffusion models generate images by explicitly modeling the noise and using a probabilistic framework to remove it step by step [ho2020denoising].
@@ 22.1.1 Forward and Reverse Diffusion Process
The forward and reverse diffusion processes are at the core of how diffusion models generate images [yang2023diffusion].
@@
16class SimpleDiffusionModel($n$.Module):
@@
19 self.denoise_net = $n$.Sequential(
20 $n$.Conv2d(3, 64, kernel_size=3, padding=1),
21 $n$.ReLU(),
22 $n$.Conv2d(64, 64, kernel_size=3, padding=1),
23 $n$.ReLU(),
24 $n$.Conv2d(64, 3, kernel_size=3, padding=1) # Predict denoised output
@@ 22.1.2 Probabilistic Diffusion Models
Probabilistic diffusion models are built on a well-defined probabilistic framework. They explicitly model the forward process of adding noise to an image and the reverse process of denoising the image. The model learns to approximate the reverse distribution $p(x_{t-1}|x_{t})$ using neural networks, which allows it to generate high-quality images [ho2020denoising].
@@ 22.2 Applications of Diffusion Models in Image Generation
Diffusion models have been successfully applied in various image generation tasks, particularly for generating high-resolution and detailed images. Two of the most popular types of diffusion models are Denoising Diffusion Probabilistic Models (DDPM) and Latent Diffusion Models (LDM), each offering unique strengths in image generation [yang2023diffusion].
@@ 22.2.1 Denoising Diffusion Probabilistic Models (DDPM)
Denoising Diffusion Probabilistic Models (DDPM) are a class of generative models that achieve state-of-the-art results in image generation by iteratively removing noise from noisy inputs. The process is highly structured, with each time step focusing on reducing a small amount of noise, making DDPMs capable of generating high-quality images [ho2020denoising].
@@
1class DDPM($n$.Module):
@@
5 self.denoise_net = $n$.Sequential(
6 $n$.Conv2d(3, 64, kernel_size=3, padding=1),
7 $n$.ReLU(),
8 $n$.Conv2d(64, 64, kernel_size=3, padding=1),
9 $n$.ReLU(),
10 $n$.Conv2d(64, 3, kernel_size=3, padding=1) # Predict denoised output
@@ 22.2.2 Latent Diffusion Models (LDM)
  • High-quality generation: Despite operating in a lower-dimensional space, LDMs can still generate high-quality images by learning a mapping between the latent space and the pixel space [rombach2022high].
@@
1class LatentDiffusionModel($n$.Module):
@@
4 self.encoder = $n$.Conv2d(3, latent_dim, kernel_size=3, padding=1)
5 self.decoder = $n$.ConvTranspose2d(latent_dim, 3, kernel_size=3, padding=1)
@@ 22.4.1 Image Denoising
6 class DiffusionModel($n$.Module):
@@
9 self.network = $n$.Sequential(
10 $n$.Conv2d(3, 64, kernel_size=3, padding=1),
11 $n$.ReLU(inplace=True),
12 $n$.Conv2d(64, 128, kernel_size=3, padding=1),
13 $n$.ReLU(inplace=True),
14 $n$.Conv2d(128, 3, kernel_size=3, padding=1)
@@ 22.4.2 Image Inpainting
1 class InpaintingDiffusionModel($n$.Module):
@@
4 self.network = $n$.Sequential(
5 $n$.Conv2d(3, 64, kernel_size=3, padding=1),
6 $n$.ReLU(inplace=True),
7 $n$.Conv2d(64, 128, kernel_size=3, padding=1),
8 $n$.ReLU(inplace=True),
9 $n$.Conv2d(128, 3, kernel_size=3, padding=1)
@@ 22.4.3 Image Super-Resolution
1 class SuperResolutionDiffusionModel($n$.Module):
@@
4 self.network = $n$.Sequential(
5 $n$.Conv2d(3, 64, kernel_size=3, padding=1),
6 $n$.ReLU(inplace=True),
7 $n$.Conv2d(64, 128, kernel_size=3, padding=1),
8 $n$.ReLU(inplace=True),
9 $n$.ConvTranspose2d(128, 3, kernel_size=3, stride=2, padding=1, output_padding=1) # Upsample
@@
20 return $n$.functional.interpolate(image, scale_factor=scale_factor, mode=’bilinear’)
@@
Deep learning has revolutionized the field of image processing, enabling powerful methods for generating, enhancing, and understanding images. From creative image synthesis to enhancing resolution and clarity, deep learning models such as GANs, diffusion models, and others have opened new avenues for solving complex image-related tasks. This chapter explores advanced applications of deep learning in image generation, enhancement, and interpretation, providing a detailed explanation of the underlying techniques and their practical implications [jiao2019survey, chen2024deeplearningmachinelearning, ren2024deeplearningmachinelearning].
@@ 23.1.1 Deep Learning-based Image Generation
Deep learning-based image generation techniques involve training models to learn the statistical patterns in a large set of images. Once trained, these models can produce novel images that resemble the patterns seen during training. Two of the most prominent approaches for image generation are Generative Adversarial Networks (GANs) and Diffusion Models. Both methods use neural networks to generate images, but their underlying mechanisms differ significantly [porkodi2023generic].
@@
: GANs are a class of deep learning models that consist of two neural networks: a generator and a discriminator, which compete in a game-like scenario. The generator tries to create realistic images, while the discriminator evaluates the images and distinguishes between real and generated ones. Over time, the generator learns to produce highly realistic images, while the discriminator becomes better at identifying fake images. This adversarial training process leads to a generator capable of creating visually appealing and detailed images [radford2015unsupervised].
@@
5class GANGenerator($n$.Module):
@@
8 self.fc = $n$.Sequential(
9 $n$.Linear(noise_dim, 256),
10 $n$.ReLU(),
11 $n$.Linear(256, 512),
12 $n$.ReLU(),
13 $n$.Linear(512, img_dim),
14 $n$.Tanh() # Output in range [-1, 1]
@@
21class GANDiscriminator($n$.Module):
@@
24 self.fc = $n$.Sequential(
25 $n$.Linear(img_dim, 512),
26 $n$.LeakyReLU(0.2),
27 $n$.Linear(512, 256),
28 $n$.LeakyReLU(0.2),
29 $n$.Linear(256, 1),
30 $n$.Sigmoid() # Output a probability
@@
  • The generator takes a random noise vector $z$ as input and transforms it into a high-dimensional image. The output is a flattened image of size 28x28 pixels, and the Tanh activation function ensures the pixel values are in the range [1,1][-1,1].
@@
  • Image-to-Image Translation: GANs are used to convert images from one domain to another, such as turning sketches into fully colored images [yi2017dualgan].
  • Style Transfer: GANs can blend the artistic style of one image with the content of another to create new, stylized images [gatys2016image].
  • Text-to-Image Generation: GANs can generate images from text descriptions, producing realistic visuals based on the input text [liao2022text].
@@
: Diffusion models represent a different approach to image generation. Unlike GANs, which use a generator-discriminator framework, diffusion models operate by progressively transforming a simple, noisy image into a high-quality image through multiple denoising steps. These models add noise to an image in a forward process and then use a learned reverse process to remove the noise step by step, eventually recovering a clear image [jiao2019survey].
@@
9class ReverseDiffusionModel($n$.Module):
@@
12 self.denoise_net = $n$.Sequential(
13 $n$.Conv2d(3, 64, kernel_size=3, padding=1),
14 $n$.ReLU(),
15 $n$.Conv2d(64, 64, kernel_size=3, padding=1),
16 $n$.ReLU(),
17 $n$.Conv2d(64, 3, kernel_size=3, padding=1) # Output denoised image
@@
Diffusion models have become particularly popular due to their stability during training and ability to generate high-resolution images [yang2023diffusion]. They are well-suited for applications that require fine-grained control over image generation, such as: